OpenStudy (anonymous):

What is the slope of the line given by the linear equation?....4x+7y=8

OpenStudy (anonymous):

r u solving for x

OpenStudy (cgreenwade2000):

You need to put the equation in slope-intercept form: y=mx+b

OpenStudy (anonymous):

7y = 8 - 4x 7y = -4x + 8 y = (-4/7)x + 8/7 y = mx + b m is the slope so your slope is -4/7

OpenStudy (goten77):